### v1

@@ 1 Introduction
The rapid growth of Large Language Models (LLMs) has revolutionized many fields, starting from natural language processing to computer vision and artificial intelligence [1, 2]. These models excel in generating coherent, contextually relevant responses by integrating textual and visual data [3]. However, their utility is hindered by a persistent issue: visual hallucinations [4] as shown in Figure 1. These hallucinations, where models generate information misaligned with the provided visual input, undermine their reliability and trustworthiness, especially in high-stakes domains [5, 6].
@@
Modern models, such as GPT-4V [7], MiniGPT [8], frequently suffer from hallucinations while generating text conditioned on visual inputs[9]. These may manifest as invented objects or relationships and raise concerns about trusting such models in real-world application scenarios [10]. With LLMs becoming integrated into fields such as healthcare, autonomous systems, and customer service, the need to look closer at such visual hallucinations has increased [5, 11].
Research on mitigating visual hallucinations can logically fall into four major areas: data-related strategies, model-related approaches, training-related techniques, and inference-related methods [12]. In this context, the common stream in the data-related strategies is the enhancement of visual instruction data to clear inaccuracies or misleading descriptions that may not match the actual visual contents[13]. Generally, model-based methods improve either the structure or the training process of LLMs in enhancing their ability to distinguish accurate information from visual displays, and data-related strategies are computationally extensive and time-consuming [14]. Our work falls under the data-related strategy, where we improve the visual instructions to get better output from the model.
@@ 2 Literature Review
Visual hallucinations in LLMs represent a critical challenge to their reliability, which has led to a great deal of research to address the issue in recent years [12, 5]. Previous studies have attempted to benchmark visual hallucinations in multimodal large language models (MLLMs). Huang et al. [9] revealed that advanced models such as GPT-4V [7], MiniGPT [8], and LLaVA [15] frequently produce hallucinated content, such as inventing non-existent objects or relationships when processing visual input. For example, GPT-4V, LLaVA-1.5 and MiniGPT-v2 showed general accuracies of 38.3%, 29.9%, and 7.5%, respectively, on the benchmark, indicating significant room for improvement. In another work, Wu et al. introduced ‘AUTOHALLUSION’ [16] introduced the first automated benchmark generation framework for probing language priors and creating visually misleading inputs via object insertion, removal, or pairing, and as a result, it exposes common biases in Large Vision Language Models (LVLMs). However, AUTOHALLUSION serves mainly as a diagnostic tool, to deliberately trigger and record hallucinations rather than trying to prevent them directly.
@@
Visual instruction data for MLLMs often include inaccurate or misleading descriptions that do not align with the actual visual content. HalluciDoctor [17] automated the detection and elimination of these hallucinations by cross-checking visual descriptions with external models such as ChatGPT to create a cleaner and more accurate dataset. To assess the method, the authors extended the CHAIR benchmark [18] with object, relation, and attribute hallucination metrics. Their model reduced object hallucinations by 44.6% over standard LLaVA. However, it relies on external tools for rewriting, which increases computational overhead.
@@
A common observation is that MLLMs are more likely to produce hallucinations when generating longer texts. To resolve this issue, instruction tuning techniques [19] have been proposed, which involve limiting the length of instruction data. Another promising approach is contrastive learning, which aligns visual and textual representations to reduce hallucinations in generated responses. Hallucination Augmented Contrastive Learning (HACL) [20] incorporated hallucinated captions as negative examples during training to help the model differentiate between accurate and hallucinated outputs. This approach has led to substantial performance gains; for instance, MiniGPT-4 and LLaVA experienced improvements of 34.66% and 29.5%, respectively, on the MMHal-Bench benchmark [21]. While effective, HACL requires resource-intensive training and is limited to improving visual-textual alignment.
@@
Kim et al. proposed ‘CODE’ [13], which uses self-created image descriptions as a guide to help the model compare and refine its responses during the decoding process. While this approach works well, it requires generating self-descriptions and figuring out contrastive references during decoding steps that can make the process more complex and slower.
@@
Yang et al. introduced ‘Pensieve’ [22], a retrieval-based technique that compares visual characteristics of similar images in a reference database (such as COCO [23]) to identify and mitigate visual hallucinations. However, this approach heavily relies on a reference dataset, which may not always be available or comprehensive.
@@
Preference Optimization methods, like HA-DPO [24] and SeVa [25], work by refining the model’s preferences through the creation of positive and negative examples. Another significant effort, Hallucination-Targeted Direct Preference Optimization (HDPO) [26], presented hallucination-targeted preference optimization by constructing diverse examples to better align MLLMs via fine-tuning, reducing hallucinations more effectively than general preference-based methods. Despite their effectiveness, these methods require extensive fine-tuning and are resource-intensive.
@@
Recent efforts have aimed to mitigate visual hallucinations using strategies such as post hoc corrections [27, 23, 28]. VOLCANO [23] uses self-feedback to iteratively evaluate and improve its answers before determining whether to accept corrections. Woodpecker [28] uses a training-free method to correct hallucinations with object detectors, Visual Question-answering (VQA) models, and GPT-based tools, focusing on post-response correction rather than prevention. On the POPE benchmark [14], this model outperforms baseline models, achieving a 30.66% accuracy improvement over MiniGPT-4 and a 24.33% gain over mPLUG-Owl [29]. But these approaches focus entirely on post-hoc corrections, without tackling the issues present at the input level.
@@
A more recent direction explores mitigation at the decoding level via latent representation manipulation. TruthPrInt [30] proposes a novel two-stage decoding-time framework. It first learns “truthful directions” from the internal hidden states of LVLMs. Then, it guides inference using a module called ComnHallu, which aligns hallucination-related latent subspaces across different models and datasets. This improves generalization and reduces hallucination on CHAIR, POPE, and LLaVA-Bench [31]. However, it requires access to model-specific hidden states, which limits its deployment in an environment where such internal access is restricted.
@@
Previous studies have utilized benchmarks such as CHAIR, POPE, MMHal-Bench and LLaVA-Bench to detect visual hallucinations in MLLMs and LVLMs. While these benchmarks provide valuable insights, they often rely on curated datasets or require access to model internals, which may not be feasible in all scenarios. In contrast, we employ SelfCheckGPT [32], a zero-resource, black-box approach that detects hallucinations by generating multiple outputs for the same prompt and assessing consistency among them using a Natural Language Inference (NLI) model. This method does not require access to the model’s internal workings or external reference data, which aligns with our objective to find out a low-resource, scalable, efficient solution.
@@ 3.1 Dataset
The samples are sourced from the HaloQuest dataset [33], with a specific focus on the first 1000 samples identified as visually challenging. This dataset consists of a diverse set of images, each paired with ground truth questions and corresponding answers, facilitating a rigorous evaluation of model performance under complex visual conditions. The structured format includes:
@@
Table 1: Dataset samples representing each of the question types used in the paper (sourced from the Holoquest Dataset [33]).
@@ 3.2 Image Preprocessing Techniques
For noise reduction, a median filter is applied to the query image. The median filter is a nonlinear spatial filter widely used for suppressing impulse noise while preserving edges [34]. Unlike linear smoothing methods, the median filter operates based on pixel intensity ordering. Given a neighborhood window $W(x,y)$ centered at pixel $(x,y)$, the output of the filter is defined as:
@@
For edge enhancement, a Laplacian filter is used, which highlights regions of rapid intensity change by computing the second-order spatial derivative of the image [35]. The continuous form of the Laplacian operator $\nabla^{2}f$ is given by:
@@
Both the noise reduction (via median filtering) and edge enhancement (via Laplacian operator) utilize a weighted image blending strategy to refine the final image output. This approach is inspired by the classical unsharp masking technique [35], which enhances image sharpness by amplifying high-frequency components.
@@
$\text{I}(x,y)=\alpha\cdot\text{src}_{1}(x,y)+\beta\cdot\text{src}_{2}(x,y)+\gamma,$ (4)
where:
  • $\text{src}_{1}$ is the original image,
  • $\text{src}_{2}$ is a filtered version (either blurred or edge-detected),
  • $\alpha$ and $\beta$ are the blending weights,
  • $\gamma$ is an optional scalar to adjust overall brightness.
@@ 3.3.1 Model
The model used in this work to detect Hallucination is SelfCheckGPT NLI [32]. SelfCheckGPT is a straightforward model that helps check if responses from language models are accurate, even without using any extra information or databases. The basic hypothesis is that if a language model knows about a topic, its answers will usually be similar and share the same facts.
@@ 3.3.2 Calculating NLI Score
The Natural Language Inference (NLI) concerns whether a premise can lead to a hypothesis. Such relationships are typically classified as entailment, neutral, and contradiction. Existing NLI metrics have also been used to evaluate the fidelity of summary generation. Maynez et al. (2020) [36], for example, used a MNLI text entailed classification model to determine whether a given summary contains internally contradictory information or not [37]. Proceeding in In NLI-focused rubric, it is likely that NLI contradiction score could as well be applied as a SelfCheckGPT score. In SelfCheck-NLI DeBERTa-v3-large (He et al., 2023) [38] a NLI model will be used which has been trained on MNLI data with a maximalist language model practiced on different languages. For most of the NLI classifiers the input consists of the premise followed by the hypothesis, which in our case is the sampled passage $S^{n}$ and the sentence under consideration $r_{i}$ as its respective constituents. Only ‘entailment’ and ‘contradiction’ classes are used, especially
@@ 4.1 Implication of Preprocessing Module
Table 2 presents a detailed quantitative analysis of how different image preprocessing techniques—namely, edge enhancement and noise reduction affect the NLI scores across various VQA categories. This table categorizes 1000 image-question pairs from the HaloQuest dataset [33] into different cases based on whether the original, edge-enhanced, or noise-reduced version yields the lowest NLI score, which serves as a proxy for reduced hallucination and improved grounding.
@@ 5 Conclusion
Ethics approval and consent to participate Not applicable
@@
Data Availability The dataset used for this research is publicly available [33].
@@
Code Availability All codes will be made publicly available upon acceptance of the manuscript.

### v2

@@ 1 Introduction
The rapid growth of Large Language Models (LLMs) has revolutionized many fields, starting from natural language processing to computer vision and artificial intelligence [hamadi2023largelanguagemodelsmeet, Budnikov2025, ajwad2025banglaCHQ]. Their increasing scale and adaptability have facilitated the deployment across a diverse spectrum of tasks, including multimodal reasoning, embodied agents, medical assistance, mental health assessment, creative text understanding, etc., demonstrating their broad applicability in both technical and socially impactful domains [Lee2025, khan-etal-2023-banglachq, ahmed2024Depression, ridwan2023poem]. These models excel in generating coherent, contextually relevant responses by integrating textual and visual data [zhang2024vllmSurvey, ivan2024meme, Wei2024]. However, their utility is hindered by a persistent issue: visual hallucinations [Huang_2025] as shown in Figure 1. These hallucinations, where models generate information misaligned with the provided visual input, undermine their reliability and trustworthiness, especially in high-stakes domains [li2025benchmark, xu2025hallucinationinevitableinnatelimitation].
@@
Modern models, such as GPT-4V [openai2023gpt4v], MiniGPT [zhu2023minigpt], frequently suffer from hallucinations while generating text conditioned on visual inputs[huang2024visual, Mohammed2025]. These may manifest as invented objects or relationships and raise concerns about trusting such models in real-world application scenarios [miyai2024unsolvable]. With LLMs becoming integrated into fields such as healthcare, autonomous systems, and customer service, the need to look closer at such visual hallucinations has increased [li2025benchmark, Rafi2025, chen2024detectingevaluatingmedicalhallucinations].
Research on mitigating visual hallucinations can logically fall into four major areas: data-related strategies, model-related approaches, training-related techniques, and inference-related methods [bai2024hallucination]. In this context, the common stream in the data-related strategies is the enhancement of visual instruction data to clear inaccuracies or misleading descriptions that may not match the actual visual contents[kim2024code]. Generally, model-based methods improve either the structure or the training process of LLMs in enhancing their ability to distinguish accurate information from visual displays, and data-related strategies are computationally extensive and time-consuming [pope]. Our work falls under the data-related strategy, where we improve the visual instructions to get better output from the model.
@@ 2 Literature Review
Visual hallucinations in LLMs represent a critical challenge to their reliability, which has led to a great deal of research to address the issue in recent years [bai2024hallucination, li2025benchmark]. Previous studies have attempted to benchmark visual hallucinations in multimodal large language models (MLLMs). Huang et al. [huang2024visual] revealed that advanced models such as GPT-4V [openai2023gpt4v], MiniGPT [zhu2023minigpt], and LLaVA [liu2023visual] frequently produce hallucinated content, such as inventing non-existent objects or relationships when processing visual input. For example, GPT-4V, LLaVA-1.5 and MiniGPT-v2 showed general accuracies of 38.3%, 29.9%, and 7.5%, respectively, on the benchmark, indicating significant room for improvement. In another work, Wu et al. introduced ‘AUTOHALLUSION’ [wu2024autohallusion] introduced the first automated benchmark generation framework for probing language priors and creating visually misleading inputs via object insertion, removal, or pairing, and as a result, it exposes common biases in Large Vision Language Models (LVLMs). However, AUTOHALLUSION serves as a diagnostic tool, to deliberately trigger and record hallucinations rather than trying to prevent them directly.
@@
Visual instruction data for MLLMs often include inaccurate or misleading descriptions that do not align with the actual visual content. HalluciDoctor [doc123] automated the detection and elimination of these hallucinations by cross-checking visual descriptions with external models such as ChatGPT to create a cleaner and more accurate dataset. To assess the method, the authors extended the CHAIR benchmark [Chair] with object, relation, and attribute hallucination metrics. Their model reduced object hallucinations by 44.6% over standard LLaVA. However, it relies on external tools for rewriting, which increases computational overhead.
@@
A common observation is that MLLMs are more likely to produce hallucinations when generating longer texts. To resolve this issue, instruction tuning techniques [liu2023mitigating] have been proposed, which involve limiting the length of instruction data. Another promising approach is contrastive learning, which aligns visual and textual representations to reduce hallucinations in generated responses. Hallucination Augmented Contrastive Learning (HACL) [jiang2023hallucination] incorporated hallucinated captions as negative examples during training to help the model differentiate between accurate and hallucinated outputs. This approach has led to substantial performance gains; for instance, MiniGPT-4 and LLaVA experienced improvements of 34.66% and 29.5%, respectively, on the MMHal-Bench benchmark [sun2023aligning]. While effective, HACL requires resource-intensive training and is limited to improving visual-textual alignment.
@@
Kim et al. proposed ‘CODE’ [kim2024code], which uses self-created image descriptions as a guide to help the model compare and refine its responses during the decoding process. While this approach works well, it requires generating self-descriptions and figuring out contrastive references during decoding steps, which can make the process more complex and slower.
@@
Yang et al. introduced ‘Pensieve’ [yang2024pensieve], a retrieval-based technique that compares visual characteristics of similar images in a reference database (such as COCO [lee2023volcano]) to identify and mitigate visual hallucinations. However, this approach heavily relies on a reference dataset, which may not always be available or comprehensive.
@@
Preference Optimization methods, like HA-DPO [zhao2023beyond] and SeVa [tan2023beyond], work by refining the model’s preferences through the creation of positive and negative examples. Another significant effort, Hallucination-Targeted Direct Preference Optimization (HDPO) [fu2023mitigating], presented hallucination-targeted preference optimization by constructing diverse examples to better align MLLMs via fine-tuning, reducing hallucinations more effectively than general preference-based methods. Despite their effectiveness, these methods require extensive fine-tuning and are resource-intensive.
@@
Recent efforts have aimed to mitigate visual hallucinations using strategies such as post hoc corrections [zhou2023analyzing, lee2023volcano, yin2023woodpecker]. VOLCANO [lee2023volcano] uses self-feedback to iteratively evaluate and improve its answers before determining whether to accept corrections. Woodpecker [yin2023woodpecker] uses a training-free method to correct hallucinations with object detectors, Visual Question-answering (VQA) models, and GPT-based tools, focusing on post-response correction rather than prevention. On the POPE benchmark [pope], this model outperforms baseline models, achieving a 30.66% accuracy improvement over MiniGPT-4 and a 24.33% gain over mPLUG-Owl [owl]. But these approaches focus entirely on post-hoc corrections, without tackling the issues present at the input level.
@@
A more recent direction explores mitigation at the decoding level via latent representation manipulation. TruthPrInt [duan2025truthprint] proposes a novel two-stage decoding-time framework. It first learns “truthful directions” from the internal hidden states of LVLMs. Then, it guides inference using a module called ComnHallu, which aligns hallucination-related latent subspaces across different models and datasets. This improves generalization and reduces hallucination on CHAIR, POPE, and LLaVA-Bench [liu2024improved]. However, it requires access to model-specific hidden states, which limits its deployment in an environment where such internal access is restricted.
@@
Previous studies have utilized benchmarks such as CHAIR, POPE, MMHal-Bench and LLaVA-Bench to detect visual hallucinations in MLLMs and LVLMs. While these benchmarks provide valuable insights, they often rely on curated datasets or require access to model internals, which may not be feasible in all scenarios. In contrast, we employ SelfCheckGPT [selff], a zero-resource, black-box approach that detects hallucinations by generating multiple outputs for the same prompt and assessing consistency among them using a Natural Language Inference (NLI) model. This method does not require access to the model’s internal workings or external reference data, which aligns with our objective to find out a low-resource, scalable, efficient solution.
@@ 3.1 Dataset
The samples are sourced from the HaloQuest dataset [wang2024haloquest], with a specific focus on the first 1000 samples identified as visually challenging. This dataset consists of a diverse set of images, each paired with ground truth questions and corresponding answers, facilitating a rigorous evaluation of model performance under complex visual conditions. The structured format includes:
@@
Table 1: Dataset samples representing each of the question types used in the paper (sourced from the Holoquest Dataset [wang2024haloquest]).
@@ 3.2 Image Preprocessing Techniques
For noise reduction, a median filter is applied to the query image. The median filter is a nonlinear spatial filter widely used for suppressing impulse noise while preserving edges [huang1979fast]. Unlike linear smoothing methods, the median filter operates based on pixel intensity ordering. Given a neighborhood window $W(x,y)$ centered at pixel $(x,y)$, the output of the filter is defined as:
@@
For edge enhancement, a Laplacian filter is used, which highlights regions of rapid intensity change by computing the second-order spatial derivative of the image [gonzalez2009digital]. The continuous form of the Laplacian operator $\nabla^{2}f$ is given by:
@@
Both the noise reduction (via median filtering) and edge enhancement (via Laplacian operator) utilize a weighted image blending strategy to refine the final image output. This approach is inspired by the classical unsharp masking technique [gonzalez2009digital], which enhances image sharpness by amplifying high-frequency components.
@@
$\textit{I}(x,y)=\alpha\cdot\text{src}_{1}(x,y)+\beta\cdot\text{src}_{2}(x,y)+\gamma,$ (4)
Where $\textit{I}(x,y)$ is the resulting pixel value at location $(x,y)$ in the image, $\text{src}_{1}$ is the original image, $\text{src}_{2}$ is a filtered version of that image (such as a blurred or edge-detected variant), $\alpha$ and $\beta$ are blending weights controlling the contribution of each source, and $\gamma$ is an optional scalar added to adjust the overall brightness of the output.
@@ 3.3.1 Model
The model used in this work to detect Hallucination is SelfCheckGPT NLI [selff]. SelfCheckGPT is a straightforward model that helps check if responses from language models are accurate, even without using any extra information or databases. The basic hypothesis is that if a language model knows about a topic, its answers will usually be similar and share the same facts.
@@ 3.3.2 Calculating NLI Score
The Natural Language Inference (NLI) concerns whether a premise can lead to a hypothesis. Such relationships are typically classified as entailment, neutral, and contradiction. Existing NLI metrics have also been used to evaluate the fidelity of summary generation. Maynez et al. (2020) [maynez-etal-2020-faithfulness], for example, used a MNLI text entailed classification model to determine whether a given summary contains internally contradictory information or not [williams-etal-2018-broad]. Proceeding in In NLI-focused rubric, it is likely that NLI contradiction score could as well be applied as a SelfCheckGPT score. In SelfCheck-NLI DeBERTa-v3-large (He et al., 2023) [he2023debertav] a NLI model will be used which has been trained on MNLI data with a maximalist language model practiced on different languages. For most of the NLI classifiers the input consists of the premise followed by the hypothesis, which in our case is the sampled passage $S^{n}$ and the sentence under consideration $r_{i}$ as its respective constituents. Only ‘entailment’ and ‘contradiction’ classes are used, especially
@@ 4.1 Implication of Preprocessing Module
Table 2 presents a detailed quantitative analysis of how different image preprocessing techniques—namely, edge enhancement and noise reduction affect the NLI scores across various VQA categories. This table categorizes 1000 image-question pairs from the HaloQuest dataset [wang2024haloquest] into different cases based on whether the original, edge-enhanced, or noise-reduced version yields the lowest NLI score, which serves as a proxy for reduced hallucination and improved grounding.
@@ 5 Conclusion
Availability of data and material The dataset used for this research is publicly available [wang2024haloquest].
[…]
Competing interests The authors declare that they have no competing interests.
@@
Authors’ contributions Nokimul Hasan Arif: Conceptualization, Methodology, Writing - original draft. Shadman Rabby: Investigation, Implementation, Visualization, Writing - original draft. Md Hefzul Hossain Papon: Formal analysis, Validation, Visualization, Writing - original draft. Sabbir Ahmed: Supervision, Project administration, Writing – review & editing, Final approval of the manuscript.
@@
Ethics approval and consent to participate Not applicable
[…]
Consent for publication Not applicable.
[…]
Acknowledgements The authors would like to thank the maintainers of the HaloQuest dataset for making their benchmark publicly available, which was instrumental to this study. We also acknowledge the support of our respective institutions in facilitating this research.